Admission to a baseball game is $4.50 for general and $5.00 for reserved seats. The receipts were $5921.00 for 1274 paid admission. How many of each ticket were sold?
how do I set up the equation and solve it?

Answer by stanbon(75887)

Admission to a baseball game is $4.50 for general and $5.00 for reserved seats. The receipts were $5921.00 for 1274 paid admission. How many of each ticket were sold?
----
# of tickets::::: g + r = 1274 tickets
value of tick:; 4.5g+5r = 5921 dollars
---
Multiply thru the 1st eq. by 50
Multiply thru the 2nd eq. by 10
-----------
50g + 50r = 50*1274
45g+50r = 59210
-----------------------
Subtract and solve for "g":
5g = 4490
g = 898 (# of general admission tickets sold)
-----
Solve for "r:
g + r = 1274
898 + r = 1274
r = 376 (# of reserve tickets sold)
======================================
Cheers,
Stan H.
==============